OpenStudy (itz_sid):

The region bounded by the given curves is rotated about the specified axis. Find the volume V of the resulting solid by any method.

OpenStudy (itz_sid):

y = −x^2 + 8x − 15, y = 0; about the x-axis Our teacher prefers us to use shells. I don't know how to solve for X :/

OpenStudy (jackthegreatest):

factor

OpenStudy (jackthegreatest):

y = (-x +3)(x-5)

OpenStudy (jackthegreatest):

split it up -x = -3 x = 5

OpenStudy (jackthegreatest):

(x = 3 or 5)

OpenStudy (jackthegreatest):

@iTz_Sid

OpenStudy (otherworldly):

its not 5 or 3.....

OpenStudy (jackthegreatest):

ok and

OpenStudy (otherworldly):

i tried it out it doesn't work

OpenStudy (jackthegreatest):

how so?

OpenStudy (jackthegreatest):

when you plug the numbers in, they check out

OpenStudy (otherworldly):

y = 0 if u put x = 3 or 5 the equation isn't 0

OpenStudy (jackthegreatest):

dont forget, it is - x^2

OpenStudy (otherworldly):

ok if it were x = - 5 it will end up - 30

OpenStudy (jackthegreatest):



OpenStudy (jackthegreatest):

but i said 3 and 5, not negative

OpenStudy (otherworldly):

i know but if u were to put the negative on the other side it would be x = -3 or -5 it means the same things as -x = 3 or 5

OpenStudy (otherworldly):

and u r correct it is -3 i only tried out -5 x = -3

OpenStudy (otherworldly):

or as u say it -x = 3

OpenStudy (jackthegreatest):

i said -x = -3

OpenStudy (jackthegreatest):

whatever, the answer is 3 and 5

OpenStudy (otherworldly):

there can't be two answers x can only be one number

OpenStudy (itz_sid):

I meant find the equation for X As in X=.....

Nnesha (nnesha):

complete the square ??hmm

OpenStudy (jackthegreatest):

hmm thats a completely different game

OpenStudy (jackthegreatest):

@OtherWorldly all that arguing... for nothing

OpenStudy (otherworldly):

yup XD

Nnesha (nnesha):

i would use disk method easy pzy

zepdrix (zepdrix):

You don't know how to solve for x? Come on Sid :d get your head in the game! Complete the square.\[\large\rm y=-(x^2-8x)-15\]Half of 8 is 4, 4 squared is 16, so I guess that completes the square for us,\[\large\rm y=-(x^2-8x+16)-15+16\]Subtract 1 to the other side,\[\large\rm y-1=-(x^2-8x+16)\]Rewrite as perfect square, multiply by negative 1,\[\large\rm 1-y=(x-4)^2\]Square root,\[\large\rm \sqrt{1-y}=x-4\]Add 4,\[\large\rm 4+\sqrt{1-y}=x\]

OpenStudy (itz_sid):

Ooooooh. xD

OpenStudy (mathmale):

Note that y = −x^2 + 8x − 15 has the graph of an inverted parabola. That is, the parabola opens down. You should find the x-coord. of the points where the graph intersects the x-axis. Draw the entire solid. Then it may become easier for you to visualize how you might use the shell method.

Nnesha (nnesha):

no luck.. i guess we still have to write the function in terms of y for disk method right ?

zepdrix (zepdrix):

In terms of y for disks? :o Hmm I don't think so. I can see that my drawing was a little off hehe. I just knew it was some kind of parabola opening downward XD